OpenStudy (anonymous):

the function y=-.03(x-14)^2+6 models the jump of a red kangaroo where x is the horizontal distance in feet and y is the corresponding height in feet. What is the kangaroos maximum height? How long is the kangaroo's jump?

OpenStudy (campbell_st):

well you'll need to find the 1st derivative set it equal to zero and solve for x. this will be the point where the maximum height occurs with respect to the horizontal distance. take the x value and substitute it into the original equation to find the max height. hope this helps.

OpenStudy (campbell_st):

to find the length of the jump... set the original equation to zero and solve for x. so \[0=-0.3(x - 14)^2 + 6\] which becomes \[-6 = -0.3(x -14)^2\] just keep doing the opposites to solve for x.